### v1

# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

Abstract
Large Language Models (LLMs) offer remarkable capabilities in code generation, natural language processing, and domain-specific reasoning. Their potential in aiding quantum software development remains underexplored, particularly for the PennyLane framework—a leading platform for hybrid quantum-classical computing. To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific code samples of quantum circuits and their contextual descriptions, specifically curated to train/fine-tune LLM-based quantum code assistance. Our key contributions are threefold: (1) the automatic creation and open-source release of a comprehensive PennyLane dataset leveraging quantum computing textbooks, official documentation, and open-source repositories; (2) the development of a systematic methodology for data refinement, annotation, and formatting to optimize LLM training efficiency; and (3) a thorough evaluation, based on a Retrieval-Augmented Generation (RAG) framework, demonstrating the effectiveness of our dataset in streamlining PennyLane code generation and improving quantum development workflows. Compared to existing efforts that predominantly focus on Qiskit, our dataset significantly broadens the spectrum of quantum frameworks covered in AI-driven code assistance. By bridging this gap and providing reproducible dataset-creation methodologies, we aim to advance the field of AI-assisted quantum programming, making quantum computing more accessible to both newcomers and experienced developers.
Index Terms:
Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), Code Completion and Assistance, Dataset Curation and Annotation, Evaluation Framework
I Introduction
Quantum computing has emerged as a transformative technology, with the potential to solve complex problems that are infeasible for classical systems [1, 2]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [3]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [4, 5, 6, 7, 8].
Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks, hindering widespread adoption [9, 10]. PennyLane is a leading open-source python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [11]. However, unlike the Qiskit framework [12], which benefits from dedicated AI-driven code assistants [13], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
In parallel, the field of Natural Language Processing (NLP) has witnessed a surge in the capabilities of LLMs, which excel at reasoning, contextual understanding, and automated code generation [14, 15, 16]. These models have been successfully leveraged in classical programming [17, 18], improving developer efficiency and reducing syntactical and logical errors. However, their application in quantum programming remains largely underexplored, particularly for PennyLane. This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for its refinement, annotation, formatting, and evaluation to facilitate LLM training.
State-of-the-Art and Limitations: While several studies have explored LLM-driven code generation for classical programming tasks [19, 20], research on quantum-specific datasets remains scarce. The most notable effort in this domain is the Qiskit Code Assistant [13], which leverages a Qiskit-specific dataset for AI-assisted quantum code generation. However, its hardware-centric approach does not extend to PennyLane, which specializes in variational quantum circuits and quantum machine learning applications [21].
The primary challenge in developing a PennyLane code assistant lies in data availability. Unlike classical programming languages, where vast repositories of well-documented code exist, quantum computing resources are fragmented across research papers, user forums, open-source repositories, and official documentation [22]. Moreover, raw quantum code samples often lack contextual annotations, making it difficult for LLMs to learn meaningful associations between quantum operations and their intended applications.
Our Approach and Contributions: In this work, we introduce PennyLang to address the gap in LLM-assisted quantum code generation for pennylane through the following novel contributions. An overview of our contributions is presented in Fig. 1.
Figure 1: Overview of novel contributions in this work.
  1. 1.
    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, two quantum computing textbooks, and the official PennyLane documentation. Each sample is meticulously annotated with explanatory comments and contextual descriptions, facilitating more effective LLM training for quantum programming.
  2. 2.
    Systematic Data Refinement and Annotation: We develop a robust methodology for dataset cleaning, eliminating non-PennyLane snippets and enriching the samples with structured metadata. To further enhance usability, we employ the GPT-4o API to convert code samples into an instruction-query format, providing clear guidance for model training in a structured, task-oriented manner.
  3. 3.
    RAG-Based Evaluation Framework: We use a Retrieval-Augmented Generation (RAG) evaluation framework to assess the dataset’s impact on model performance. The framework compares outputs from three models—GPT-4o mini, Claude 3.5, and Qwen 7B—under two conditions: (i) with access to the dataset as retrieval-augmented context and (ii) without dataset augmentation. The generated outputs are validated and scored based on functionality, syntax correctness, and modularity.
Summary of Key Results: RAG-enhanced quantum code generation significantly improved performance across all models. GPT-4o Mini showed an 11.67% gain in accuracy and modularity, Claude 3.5 Sonnet improved by 7.6%, enhancing coherence, while Qwen 2.5 saw the highest boost at 14.38%, benefiting from retrieved context. These results confirm RAG’s effectiveness in optimizing PennyLane quantum code generation by improving functionality, syntax, and adaptability.
Paper Organization: Sec. II provides background information and reviews related work. Sec. III outlines our methodology for dataset collection, cleaning, and annotation along with the tokenization and formatting pipeline used for preprocessing. Sec. IV presents experimental results, including an evaluation of our dataset using a RAG-based framework. Sec. V presents some case studies Finally, Sec. VI concludes the paper with a summary of findings and future research directions.
II Background and Related Work
TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training
Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Textbook Dataset [23] Qiskit Qiskit.org Textbook Multiple structured examples Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [24] Qiskit IBM Quantum Experience Derived from past challenges Instruction-Response Format JSON/Notebook Qiskit Code Repository [25] Qiskit GitHub (Qiskit AI repos) Thousands (Scattered) Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [26] Qiskit Kaggle (Possible) 309 (Limited samples) Pretrained Benchmark Data CSV/JSON QDataSet [27] Qiskit GitHub 52 (raw datasets) Quantum ML Data Various QCircuitNet [28] Qiskit Research Paper 4,800 Samples Quantum Algorithm Design Benchmark Python/QASM PennyLane Documentation [29] PennyLane PennyLane.ai Docs Various examples Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [30] PennyLane GitHub Open-Source Scattered code samples Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Instruction-Response Dataset Curated and Structured Dataset JSON file
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [31] and automated code generation [32], their application in quantum software development remains under-explored. Existing studies on AI-driven code assistants primarily focus on classical programming languages such as Python, Java, and C++, or quantum computing frameworks like Qiskit [13]. However, the PennyLane framework [11], widely used for hybrid quantum-classical computing, lacks dedicated research efforts in AI-assisted quantum programming.
This section presents an in-depth review of related work in three key areas:
  1. 1.
    LLMs for Code Generation in Classical Programming
  2. 2.
    LLMs for Quantum Programming
  3. 3.
    Datasets for Training LLMs in Quantum Code Generation (summarized in Table I)
We then summarize existing contributions and highlight our novel approach in developing a PennyLane-centric dataset to advance AI-driven quantum programming, as shown in Table II.
II-A LLMs for Code Generation in Classical Programming
Transformer-based models have significantly impacted AI-driven programming, automating code generation, completion, and debugging. Key contributions include:
  • Codex (OpenAI) [17] introduced a language model trained on public code repositories, enabling natural language-to-code translation.
  • CodeBERT [33] leveraged bidirectional transformers to enhance source code understanding.
  • InCoder [20] developed a multi-turn interactive model for code infilling and synthesis.
  • AlphaCode [16] showed competitive programming-level LLM capabilities.
Additionally, benchmarking studies have evaluated LLM performance on programming tasks:
  • HumanEval [15] proposed an execution-based evaluation for Python code generation.
  • MBPP [34] introduced a multi-language benchmark covering correctness and execution constraints.
  • CodeXGLUE [35] aggregated multiple datasets to benchmark code-related tasks.
These studies highlight the necessity of well-structured datasets to enhance LLM code generation performance.
II-B LLMs for Quantum Programming
Existing research on AI-powered quantum code assistants focuses primarily on the Qiskit framework:
  • Qiskit Code Assistant (IBM Research) [13] assists users in debugging and optimizing Qiskit code.
  • Qiuchi et al. [36] adapted BERT-based models to improve quantum code understanding.
  • CodeLlama-Quantum (Meta AI) [18] extended CodeLlama for Qiskit-based applications.
Despite these advancements, challenges persist:
  1. 1.
    Limited Training Data: Quantum programming lacks large-scale, high-quality datasets.
  2. 2.
    Complexity of Quantum Circuits: Unlike classical code, quantum programs require understanding quantum mechanics.
  3. 3.
    Hybrid Workflows: PennyLane integrates classical and quantum processing, requiring LLMs to learn hybrid paradigms.
These challenges underscore the need for PennyLane-specific datasets for LLM training, as none currently exist, reinforcing the importance of our research.
Figure 2: Methodology for PennyLang quantum code generation and evaluation. Data is collected from GitHub, quantum books, and PennyLane docs, then refined, tokenized, and formatted for LLMs. The curated dataset is processed by GPT-4o for input-output conversion. User queries are handled via a retrieval-augmented generation (RAG) framework using LangChain and Chroma to fetch relevant code. Responses from GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B-Instruct-Turbo are compared against non-RAG outputs and evaluated on functionality, syntax, and modularity to assess RAG’s impact.
Our contributions advance AI-assisted quantum programming by enabling LLMs to provide high-quality, domain-specific assistance in PennyLane.
TABLE II: Comparison of Related Work and Our Contributions
Research Area Existing Works Our Contributions
LLMs for Code Generation Codex [17], CodeBERT [33] PennyLane-specific dataset
LLMs for Quantum Programming Qiskit Assistant [13], Qiuchi et al. [36] Hybrid quantum-classical dataset
Quantum Code Datasets QCircuitNet [28], QDataset  [28] First PennyLane dataset
III PennyLang Framework
III-A Methodology for Data Collection and Refinement
Fig. 2 illustrates the methodology employed to construct a high-quality dataset of PennyLane code samples, ensuring that each snippet is genuinely relevant, well-structured, and accurately annotated. Our data collection strategy focused on extracting, filtering, and refining code snippets from multiple sources, including GitHub repositories, quantum computing books, and official PennyLane documentation.
III-A1 Data Collection from GitHub Repositories
To systematically gather PennyLane-specific code, we utilized the GitHub API to search for publicly accessible repositories with permissive open-source licenses (e.g., MIT, Apache 2.0, BSD). The selection criteria required that the term "PennyLane" be present in either the repository name, description, or README file.
To maintain dataset quality and avoid redundancy, we followed these constraints:
  • Only repositories from the main branch at the latest commit were considered.
  • Forked repositories were systematically excluded.
  • Only .py files (Python scripts) were extracted for further analysis.
  • Non-PennyLane files were filtered by checking for explicit import pennylane as qml statements.
  • Large, autogenerated files (e.g., test logs, package files) were excluded.
Since PennyLane maintains backward compatibility and rarely deprecates features [11], we did not impose a time constraint on data collection, ensuring a broad dataset scope. Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code.
This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [30].
III-A2 Extracting Quantum Code from Books
In addition to GitHub, we extracted PennyLane-related code snippets from two quantum computing books:
  • “Quantum Machine Learning: An Applied Approach” [21]
  • “A Practical Guide to Quantum Machine Learning and Quantum Optimization” [37]
These books provide theoretical foundations and applied examples of PennyLane-based quantum machine learning (QML) and variational quantum algorithms (VQAs). The extraction process involved:
  1. 1.
    Identifying all PennyLane-related code snippets.
  2. 2.
    Diversity Assurance: The dataset includes a broad spectrum of PennyLane use cases, from basic gate operations to advanced quantum optimization problems.
  3. 3.
    Extracting only the relevant code while discarding unrelated sections.
  4. 4.
    Manually verification of each snippet with explanations from the books, converting descriptive text into structured comments.
Unlike structured repositories, book content lacked a consistent format for where explanations were placed (before or after code). Thus, manual review was necessary to ensure that annotations correctly corresponded to each snippet.
III-A3 Scraping Code from PennyLane Documentation
The official PennyLane documentation [29] serves as a comprehensive reference for quantum programming, covering:
  • Quantum circuits and variational algorithms.
  • Quantum operators and templates.
  • Hybrid quantum-classical optimization.
  • Quantum chemistry applications.
TABLE III: Tokenized Input and Attention Masking
Tokenized Input Attention Mask
"<pad>" 0
"<pad>" 0
"def" 1
":" 1
"circuit" 1
Tokenized Input Attention Mask
"(" 1
")" 1
"return" 1
"qml.expval" 1
To further enhance our dataset, we systematically scraped all code snippets from the official PennyLane website. This included examples embedded in tutorials and API references.
As with book sources, documentation explanations were manually converted into comments, ensuring that each code sample contained self-contained, contextualized information. This practice enhances LLM interpretability, as models trained on such data can generate both code and explanations.
III-A4 Dataset Summary and Statistics
The final dataset contains a total of 3,347 PennyLane code samples, as shown in Table IV.
TABLE IV: Dataset Composition
Source Samples Collected
Unofficial GitHub repositories 1,952
GitHub (PennyLaneAI repository + others) 1,321
Official PennyLane Documentation 53
Quantum Computing Books 21
Total Dataset Size 3,347
Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for fine-tuning large language models (LLMs) for PennyLane-based quantum programming.
III-A5 Ensuring Data Integrity and Diversity
To maintain dataset integrity, we applied multiple quality control measures:
  1. 1.
    Duplicate Removal: Identical code samples from different sources were eliminated.
  2. 2.
    Code Formatting: Samples were formatted consistently to follow PEP 8 style guidelines [38].
  3. 3.
    Manual Review: A human-in-the-loop verification was conducted to assess annotation accuracy.
These efforts ensure that the dataset is rich, well-structured, and optimized for AI model training.
III-A6 Implications for AI-Driven PennyLane Code Generation
The collected dataset significantly enhances AI-driven PennyLane code assistance in multiple ways:
  • Improving Model Performance: High-quality annotations ensure that models understand both syntax and context.
  • Facilitating Explainable AI: Models trained on our dataset can generate explanatory comments alongside code.
This dataset serves as a foundation for future research in AI-assisted quantum programming for PennyLane.
TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework
Component Function Description
DirectoryLoader Document Loader Loads PennyLane dataset samples as text files from a directory.
Chroma Vector Database Stores embeddings for retrieval using similarity search.
OpenAIEmbeddings Embedding Model Converts text into vector embeddings for indexing and retrieval.
ChatPromptTemplate Prompt Formatting Defines structured prompts for AI models to ensure consistency.
create_retrieval_chain Retrieval Pipeline Creates a pipeline that integrates stored embeddings for improved responses.
ChatOpenAI OpenAI LLM Interface Connects to OpenAI models (GPT-4o-mini) for response generation.
ChatAnthropic Anthropic LLM Interface Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses.
TABLE VI: LLMs Used in the RAG-Based Evaluation Framework
Model Parameters Architecture Provider Usage in Pipeline
Qwen2.5-7B-Instruct-Turbo 7B Causal Decoder-Only Transformer Alibaba Cloud Generates PennyLane code
GPT-4o Mini Not disclosed Causal Decoder-Only Transformer OpenAI Evaluates and scores generated responses
Claude 3.5 Sonnet Not disclosed Causal Decoder-Only Transformer Anthropic Alternative model for response generation
III-B Tokenization and Formatting for LLM Training
This section details the preprocessing pipeline for preparing our dataset for (LLM) training/fine-tuning, focusing on tokenization, padding strategies, and attention masking to optimize input handling within the transformer architecture.
III-B1 Handling Variable-Length Code Samples
The collected PennyLane code samples exhibit varying lengths after tokenization. Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$. Following standard practice, left padding was applied to align shorter samples to $N$ tokens, ensuring batch processing efficiency and minimizing computational overhead.
III-B2 Padding Strategies and Tokenizer Variations
Padding strategies differ across tokenizers:
  • Token-Specific Padding: Some tokenizers have a dedicated padding token, automatically appended to shorter sequences. For example, Hugging Face’s Tokenizer uses "<pad>" by default [39].
  • Manually Defined Padding: If a tokenizer lacks an explicit padding token, a placeholder token (e.g., "0" or "<mask>") must be manually assigned.
  • Left vs. Right Padding: Left padding is commonly used in autoregressive models [40], ensuring that padding tokens do not interfere with causal attention mechanisms.
In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, and Claude 3.5 Sonnet. This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining autoregressive consistency.
III-B3 Applying Attention Masking
To improve transformer efficiency, an attention mask [41] was applied post-padding. This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data. Table III represents the attention mask matrix applied to a batch of tokenized samples. Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
  • Attention Mask Structure: A binary mask is assigned to each token:
    • 1 → Meaningful tokens to be processed by the model.
    • 0 → Padding tokens ignored by the transformer.
  • Impact on Training: The self-attention mechanism [40] ensures that padding tokens do not contribute to gradients, thereby reducing unnecessary computations and preventing misleading dependencies.
III-C Methodology for Dataset Evaluation
III-C1 Retrieval-Augmented Generation (RAG) Framework
To enhance LLM-based quantum code generation, we employ a Retrieval-Augmented Generation (RAG) methodology that integrates retrieval-based context with language model inference. This hybrid approach significantly improves:
  • Accuracy: By grounding responses in retrieved dataset samples, reducing model hallucination.
  • Adherence to PennyLane’s best practices: Ensuring generated quantum code aligns with official standards.
  • Response Quality: Structured retrieval enriches contextual understanding, improving coherence and correctness.
III-C2 RAG Implementation using LangChain
LangChain is utilized to construct a modular and efficient RAG pipeline. The framework consists of the following core components:
  1. 1.
    Dataset Loading: PennyLane dataset samples are extracted from curated dataset structured in intruction-response format.
  2. 2.
    Embedding Generation: Each sample is tokenized and converted into high-dimensional vector embeddings using OpenAIEmbeddings.
  3. 3.
    Vector Storage and Retrieval: The embeddings are stored in a Chroma vector database for fast retrieval.
  4. 4.
    Prompting and Response Generation:
    • When a query is submitted, Maximum Marginal Relevance (MMR) search fetches the most relevant samples in RAG-augmented input.
    • Retrieved context is formatted into a structured prompt.
    • The LLM model generates responses using both RAG-augmented and non-RAG prompts.
  5. 5.
    Evaluation: The responses are compared using a second LLM (GPT-4o Mini), which scores them based on:
    • Functionality: Logical correctness of the quantum circuit.
    • Syntax: Validity of the PennyLane code.
    • Modularity: Adherence to best practices in structuring quantum workflows.
Figure 3: Performance of GPT 4o mini, Claude, and Qwen With and Without RAG. Mean scores (±1 SD) over 29 trials for GPT 4o mini, Claude, and Qwen under both conditions (blue: With RAG; orange: Without RAG), with individual scores shown as swarmplot markers.
Figure 3 presents the results of our RAG-based evaluation methodology, demonstrating a significant performance improvement with retrieval-augmented responses. Each model was tested with 30 code samples, and responses were evaluated using GPT-4o, which graded them based on syntax, functionality, and modularity. The results indicate that the RAG-enhanced approach consistently outperforms the base LLM, yielding higher overall scores.
III-C3 LLM Models and Configuration
The models used in this framework differ in size, architecture, and inference capabilities. Table VI provides a comprehensive comparison of key characteristics including parameter count, context window size, and quantum-specific training data percentage.
Figure 4: Distribution of instruction and response lengths with their correlation. The left plot shows the normal distribution of instruction lengths, the middle plot highlights the right-skewed distribution of response lengths, and the right scatter plot illustrates the relationship between instruction and response length.
III-C4 Prompt Design for RAG and Non-RAG Evaluation
To ensure fair evaluation, we construct two different prompts—one for RAG-augmented inference and another for direct model inference without retrieval.
  • Non-RAG Prompt:
    Base LLM Prompt (Without RAG) "You are an advanced assistant for quantum computing tasks using PennyLane. Provide a response to the user’s question. Make sure to follow best practices and optimize for efficiency. If you are unsure, state that clearly."
  • RAG-Enhanced Prompt:
    RAG-Enhanced Prompt with Context Retrieval "You are an advanced assistant for quantum computing tasks using PennyLane. Use the following pieces of retrieved context to answer the question. Use multiple retrieved code snippets whenever necessary to improve accuracy. If a single snippet does not provide enough information, retrieve and combine relevant parts from multiple sources. Make sure to follow best practices and optimize for efficiency. If you don’t know the answer, say that you don’t know." "{context}"
III-C5 Evaluation Criteria and Scoring
To assess the effectiveness of RAG-based retrieval, generated responses are evaluated using GPT-4o with the following scoring prompt:
Evaluation Prompt for GPT-4o You are an AI assistant tasked with evaluating two different answers to the same question. Question: {question} Answer with RAG: {rag_answer} Answer without RAG: {non_rag_answer} Please provide a score from 1 to 30 for each answer based on: - Functionality (correctness of quantum operations) - Syntax (validity of PennyLane code) - Modularity (code organization and efficiency) Justify your scoring.
III-C6 Findings and Impact of RAG Integration
After evaluating multiple quantum programming queries using the RAG vs. Non-RAG framework, we observe: Based on the evaluation results, RAG-enhanced responses consistently outperformed non-RAG responses across all tested models—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5. On average, RAG integration improved response scores by 11.67% in GPT-4o Mini, 7.69% in Claude 3.5, and 14.38% in Qwen 2.5. The improvement was most noticeable in functionality and syntax, as retrieved context helped prevent incorrect quantum operations and improved PennyLane-specific syntax accuracy. Claude 3.5 Sonnet demonstrated higher coherence and structured responses, while GPT-4o Mini produced concise yet effective outputs. Qwen 2.5 benefited the most from RAG, showing a 14.38% increase in performance, likely due to its stronger reliance on retrieved context for accurate quantum code generation. These results highlight the effectiveness of retrieval-augmented methods in enhancing quantum code generation accuracy and adherence to best practices.
IV Statistical Evaluation of the Dataset
A comprehensive analysis of the dataset was conducted to assess the distribution of instruction lengths, response lengths, response variations across feature groups, and their correlation with quantum functionalities. This section provides insights into the dataset’s structure, response variability, and its implications for training LLMs.
IV-A Response Lengths Across Feature Groups
Figure 5 provides a box plot representation of response lengths across different feature groups.
  • Response lengths remain relatively stable across feature groups, suggesting uniformity in generated outputs.
  • Outliers appear in every feature category, indicating occasional long responses, particularly in core operations, measurement, and mathematical functions.
  • The “Other” category has the shortest responses, likely due to simpler queries requiring minimal output.
Figure 5: Box plot of response lengths across different feature groups. Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions
IV-B Instruction and Response Length Distributions
Figure 4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset. Instruction lengths, as shown in the left plot, follow a normal distribution with a peak around 35 words. The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare. In contrast, response lengths exhibit a right-skewed distribution, as shown in the middle plot. Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs. The instruction-response correlation, illustrated in the rightmost scatter plot, reveals a positive correlation between instruction and response length. While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
IV-C Feature Group Usage Across Quantum Domains
The relationship between feature groups and quantum computing domains, such as machine learning, hardware, chemistry, and optimization, is illustrated in Figure 6. Quantum Machine Learning relies predominantly on quantum gates and operators, alongside error handling techniques, reflecting its dependence on structured quantum transformations and noise mitigation. In contrast, Quantum Chemistry and Optimization exhibit relatively lower usage of mathematical functions, suggesting that these domains rely on alternative computational approaches rather than extensive arithmetic operations. Quantum Hardware applications, however, demonstrate a balanced distribution across various feature groups, indicating a broader, more integrated usage of quantum functionalities across different operational layers.
Figure 6: Feature Group Usage Across Domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories, with color intensity reflecting count.
IV-D Feature Category Association with Quantum Functions
To analyze how different quantum functionalities relate to feature categories, we examine the confusion matrix in Figure 7.
Figure 7: Confusion matrix showing the relationship between newly defined semantic categories and various PennyLane quantum features. The color intensity represents the frequency of occurrences, where darker shades indicate higher associations between a given category (rows) and a specific quantum feature (columns). This visualization provides insights into how different quantum functionalities are distributed across conceptual categories, aiding dataset structuring and LLM-based learning tasks.
Quantum gates and operators exhibit the strongest associations with multiple features, particularly with commonly used functions such as qml.PauliZ, qml.RX, and qml.RY, which are integral to quantum circuit construction. Core operations and measurement functions also display high correlations, reflecting their frequent application in quantum circuit execution and result interpretation. Meanwhile, error handling and mathematical functions show relatively lower but still significant interactions with select quantum features, indicating their importance in specific computational scenarios rather than across all use cases.
IV-E Insights and Implications
The findings from Figures 4, 5, 7, and 6 offer several key insights:
  • Balanced Instruction Lengths: The normal distribution of instruction lengths ensures the dataset is well-structured for model training.
  • Response Length Variability: Most responses are concise, but outliers indicate that certain quantum features demand more detailed explanations.
  • Feature-Specific Trends: Quantum gates and measurement functions dominate dataset interactions, aligning with core PennyLane applications.
  • Domain-Specific Usage: Different quantum fields rely on distinct feature sets, highlighting potential customization opportunities for specialized LLM training.
V Case Studies
The case studies compare base LLM-generated PennyLane code with RAG-enhanced implementations, showcasing notable improvements in functionality, syntax, and modularity across different quantum programming tasks.
In the first task, implementing a 1-qubit parameterized RX gate, the base model introduced redundant imports, suboptimal measurement choices, and missing hardware-specific optimizations. The RAG-optimized implementation addressed these issues by incorporating hardware-aware device initialization, optimized measurement practices, and type annotations, achieving a perfect 10/10 functionality score.
The second task, developing an RX gate circuit with optional parameter binding, revealed static parameter handling and the absence of deferred execution in the base LLM output. The RAG-enhanced version introduced flexible parameter binding, probabilistic measurement strategies, and deferred execution patterns, improving adaptability and modularity.
The third case study, constructing a variational quantum classifier using StronglyEntanglingLayers, further demonstrated the superiority of RAG-based generation. The base LLM solution lacked adaptive training integration, efficient optimization, and TensorFlow-Keras support, requiring manual gradient updates. In contrast, the RAG-optimized model seamlessly integrated TensorFlow, Keras, and the Adam optimizer, significantly enhancing training efficiency and parameter optimization.
Overall, across all tasks, RAG-assisted code generation consistently outperformed the base LLM, producing more structured, modular, and functionally superior PennyLane code. This highlights the effectiveness of retrieval-augmented approaches in enhancing quantum programming workflows.
Task: Implement 1-qubit parameterized RX gate using PennyLane
Base LLM Implementation (Without RAG) Limitations Identified: Unnecessary numpy import Suboptimal measurement choice (expectation value vs state) Missing hardware-specific optimizations 1import pennylane as qml 2from pennylane import numpy as np # Redundant import 3 4dev = qml.device(’default.qubit’, wires=1) 5 6@qml.qnode(dev) 7def circuit(theta): 8 qml.RX(theta, wires=0) 9 return qml.expval(qml.PauliZ(0)) # Non-optimal measurement Listing 1: Non-RAG Implementation Evaluation Results: Functionality: 9/10 (Valid but limited measurement) Syntax: 9/10 (Redundant imports) Modularity: 8/10 (Missing type hints)
RAG-Enhanced Implementation Retrieved Context: Hardware-aware device initialization (sample_1137.txt) State measurement best practices (sample_214.txt) Testing methodologies (sample_23.txt) 1import pennylane as qml 2 3dev = qml.device("lightning.qubit", wires=1) # Hardware-optimized 4 5@qml.qnode(dev) 6def circuit(theta: float): # Type hinting 7 qml.RX(theta, wires=0) 8 return qml.state() # Full state measurement Listing 2: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Complete state analysis) Syntax: 10/10 (Clean imports) Modularity: 9/10 (Type annotations)
Task: Create 1-qubit RX gate circuit with optional parameter binding
Base LLM Implementation (Without RAG) Limitations Identified: Static parameter handling Missing deferred parameterization Limited measurement flexibility 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 if theta: # Binary parameter check 6 qml.RX(theta, wires=0) 7 return qml.expval(qml.PauliZ(0)) # Fixed measurement Listing 3: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (No variable binding) Syntax: 9/10 (Valid but rigid) Modularity: 8/10 (No late binding support)
RAG-Enhanced Implementation Retrieved Context: Variable parameter techniques (sample_1137.txt) Probabilistic measurement strategies (sample_214.txt) Deferred execution patterns (sample_23.txt) 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 param = theta if theta else qml.variable(’theta’) 6 qml.RX(param, wires=0) 7 return qml.probs(wires=0) # Full probability measurement Listing 4: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Flexible parameter binding) Syntax: 9/10 (Clear variable handling) Modularity: 9/10 (Supports late binding)
Task: Implement a Variational Quantum Classifier with Strongly Entangling Layers
Base LLM Implementation (Without RAG) Limitations Identified: Uses a static quantum circuit with fixed parameters, lacking adaptive training integration. Lacks integration with TensorFlow/Keras, requiring a manual optimization loop. Uses Gradient Descent Optimizer instead of the Adam optimizer, which may slow convergence. 1import pennylane as qml 2import numpy as np 3from sklearn.datasets import make_classification 4from sklearn.model_selection import train_test_split 5from sklearn.preprocessing import StandardScaler 6from sklearn.metrics import accuracy_score 7 8# Generate a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Scale the features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define the quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Define the quantum circuit 21@qml.qnode(dev) 22def circuit(params, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.StronglyEntanglingLayers(params, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 5: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (Works but lacks adaptive training integration) Syntax: 9/10 (Code is valid but manual optimization is not ideal) Modularity: 8/10 (Training and inference should be better structured)
RAG-Enhanced Implementation Retrieved Context: TensorFlow-Keras integration for efficient optimization. Adam optimizer instead of manual updates for improved convergence. Improved quantum circuit design based on best practices. 1import pennylane as qml 2import numpy as np 3import tensorflow as tf 4from sklearn.datasets import make_classification 5from sklearn.model_selection import train_test_split 6from sklearn.preprocessing import StandardScaler 7 8# Create a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Standardize features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Quantum circuit with Strongly Entangling Layers 21@qml.qnode(dev) 22def quantum_circuit(weights, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.templates.StronglyEntanglingLayers(weights, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 6: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Seamless Keras integration for training) Syntax: 10/10 (Concise and efficient TensorFlow pipeline) Modularity: 9/10 (Well-structured training process)
VI Conclusion
This paper presents a dedicated PennyLane dataset and a systematic methodology for preparing quantum code snippets for LLM training. Our open-source release aims to empower researchers and developers to build advanced quantum code assistants that reduce onboarding difficulties and accelerate development in hybrid quantum-classical systems. Ongoing work includes expanding the dataset with more domain-specific samples, conducting broader evaluations with different LLMs, and exploring cross-framework synergy for universal quantum code generation.
Acknowledgment
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008, and Center for CyberSecurity (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104.