### v2

# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

Abstract
Large Language Models (LLMs) offer remarkable capabilities in code generation, natural language processing, and domain-specific reasoning. However, their application in quantum software development remains underexplored, particularly for PennyLane—a leading framework for hybrid quantum-classical computing. To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific quantum code samples and contextual descriptions, specifically curated to support LLM training and fine-tuning for quantum code assistance. Our contributions are threefold: (1) the automatic construction and open-source release of a comprehensive PennyLane dataset derived from textbooks, official documentation, and open-source repositories; (2) a structured methodology for data curation, annotation, and formatting to enhance LLM usability and relevance; and (3) a rigorous evaluation of code generation capabilities using both baseline Retrieval-Augmented Generation (RAG) and a GraphRAG-enhanced pipeline. Using the PennyLang framework, we demonstrate that GraphRAG, when applied to a GPT-4o Mini model, substantially outperforms standard prompting and baseline RAG. Accuracy improves from 20.5% (without RAG) to 58.2% with GraphRAG, showcasing its effectiveness in reducing hallucinations and improving code correctness in quantum programming tasks. Compared to prior efforts focused largely on Qiskit, our work expands LLM-based assistance to the PennyLane ecosystem, contributing practical tools and reproducible methodologies for advancing AI-assisted quantum software development.
Index Terms:
Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), GraphRAG, Code Completion and Assistance, Dataset Curation, Evaluation Framework
I Introduction
Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [1, 2, 3]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [4]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [16, 17, 18]. PennyLane is a leading open-source python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [19]. However, unlike the Qiskit framework [20, 21], which benefits from dedicated AI-driven code assistants [22], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
In parallel, the field of Natural Language Processing (NLP) has seen rapid advancements in Large Language Models (LLMs), which now excel at reasoning, contextual understanding, and automated code generation [23, 24, 25]. These models have shown success in classical programming [26, 27], improving developer productivity and reducing syntax and logic errors. However, their application in quantum programming—especially within the PennyLane ecosystem—remains underexplored. This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for refinement, annotation, and evaluation, ultimately enabling LLM-based quantum code generation and debugging.
State-of-the-Art and Limitations: While LLMs have been benchmarked extensively for classical programming [28, 29], their performance in quantum domains is less understood. Qiskit HumanEval and Qiskit Code Assistant [22] represent early attempts to extend LLM capabilities into quantum programming, but they focus primarily on Qiskit—a hardware-centric framework. PennyLane, in contrast, is designed for variational quantum algorithms and quantum machine learning [30], requiring a different kind of LLM support and training corpus.
The primary challenge in developing a PennyLane code assistant lies in data availability. Unlike classical programming languages, where vast repositories of well-documented code exist, quantum computing resources are fragmented across research papers, user forums, open-source repositories, and official documentation [31]. Moreover, raw quantum code samples often lack contextual annotations, making it difficult for LLMs to learn meaningful associations between quantum operations and their intended applications.
Our Approach and Contributions: In this work, we introduce PennyLang, a large-scale, high-quality dataset tailored for PennyLane code generation, along with a rigorous evaluation pipeline. In addition to a baseline RAG framework, we propose and test the GraphRAG framework [32], which leverages structured code relationships to improve the quality and relevance of retrieval during generation. Our contributions are summarized in Fig. 1 and detailed below:
Figure 1: Overview of novel contributions in this work.
  1. 1.
    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, quantum computing textbooks, and official PennyLane documentation. Each sample is enriched with contextual descriptions and annotations to support instruction tuning.
  2. 2.
    Systematic Data Refinement and Annotation: We present a robust methodology for dataset cleaning and metadata enrichment. Using the GPT-4o API, code snippets are transformed into instruction-query pairs, improving compatibility with LLM training pipelines.
  3. 3.
    RAG-Based Evaluation Framework: We use a Retrieval-Augmented Generation (RAG) evaluation framework to assess the dataset’s impact on model performance. The framework compares outputs from three models—GPT-4o mini, Claude 3.5, and Qwen 7B—under two conditions: (i) with access to the dataset as retrieval-augmented context and (ii) without dataset augmentation. The generated outputs are validated and scored based on functionality, syntax correctness, and modularity.
  4. 4.
    GraphRAG-Enhanced Evaluation Framework: We extend the baseline RAG framework with GraphRAG, which introduces a graph-based context retrieval mechanism for code augmentation. GraphRAG models contextual and structural relationships between quantum code samples, enabling more accurate and consistent retrievals. We evaluate this setup using GPT-4o Mini model, comparing GraphRAG, and no-RAG scenarios in a controlled setting.
Summary of Key Results: RAG-enhanced quantum code generation improves performance across all models. GPT-4o Mini showed an 11.67% gain in accuracy and modularity, Claude 3.5 Sonnet improved by 7.6%, enhancing coherence, while Qwen 2.5 saw the highest boost at 14.38%, benefiting from retrieved context. These results confirm RAG’s effectiveness in optimizing PennyLane quantum code generation by improving functionality, syntax, and adaptability. Our results from GraphRAG significantly outperforms both baseline RAG and vanilla prompting. When applied to GPT-4o Mini, GraphRAG boosts accuracy from 20.5% (no RAG) to 58.2%, marking a 37.7% absolute improvement in correct solutions across PennyLang dataset, confirming the benefit of structured retrieval in quantum programming tasks.
Paper Organization: Sec. II reviews related work and background on quantum programming and RAG. Sec. III outlines our dataset construction, annotation methodology, and the GraphRAG retrieval pipeline. Sec. IV presents experimental results comparing model performance across prompting settings. Sec. V discusses case studies and qualitative insights. Sec. VI concludes with future directions for dataset-driven quantum code assistance.
II Background and Related Work
TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training.
Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Textbook Dataset [33] Qiskit Qiskit.org Textbook Multiple structured examples Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [34] Qiskit IBM Quantum Experience Derived from past challenges Instruction-Response Format JSON/Notebook Qiskit Code Repository [35] Qiskit GitHub (Qiskit AI repos) Thousands (Scattered) Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [36] Qiskit Kaggle (Possible) 309 (Limited samples) Pretrained Benchmark Data CSV/JSON QDataSet [37] Qiskit GitHub 52 (raw datasets) Quantum ML Data Various QCircuitNet [38] Qiskit Research Paper 4,800 Samples Quantum Algorithm Design Benchmark Python/QASM PennyLane Documentation [39] PennyLane PennyLane.ai Docs Various examples Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [40] PennyLane GitHub Open-Source Scattered code samples Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Instruction-Response Dataset Curated and Structured Dataset JSON file
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [41] and automated code generation [42], their application in quantum software development remains underexplored. Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [22], with little attention to the PennyLane framework [19], despite its popularity in hybrid quantum-classical workflows.
This section presents an in-depth review of related work in three key areas:
  1. 1.
    LLMs for Code Generation in Classical Programming
  2. 2.
    LLMs for Quantum Programming
  3. 3.
    Datasets for Training LLMs in Quantum Code Generation (summarized in Table I)
We then summarize existing contributions and highlight our novel approach in developing a PennyLane-centric dataset to advance AI-driven quantum programming, as shown in Table II.
II-A LLMs for Code Generation in Classical Programming
Transformer-based models have significantly impacted AI-driven programming, automating code generation, completion, and debugging. Key contributions include:
  • Codex (OpenAI) [26] introduced a language model trained on public code repositories, enabling natural language-to-code translation.
  • CodeBERT [43] leveraged bidirectional transformers to enhance source code understanding.
  • InCoder [29] developed a multi-turn interactive model for code infilling and synthesis.
  • AlphaCode [25] showed competitive programming-level LLM capabilities.
Additionally, benchmarking studies have evaluated LLM performance on programming tasks:
  • HumanEval [24] proposed an execution-based evaluation for Python code generation.
  • MBPP [44] introduced a multi-language benchmark covering correctness and execution constraints.
  • CodeXGLUE [45] aggregated multiple datasets to benchmark code-related tasks.
These studies highlight the necessity of well-structured datasets to enhance LLM code generation performance.
II-B LLMs for Quantum Programming
Existing research on AI-powered quantum code assistants focuses primarily on the Qiskit framework:
  • Qiskit Code Assistant (IBM Research) [22] assists users in debugging and optimizing Qiskit code.
  • Qiuchi et al. [46] adapted BERT-based models to improve quantum code understanding.
  • CodeLlama-Quantum (Meta AI) [27] extended CodeLlama for Qiskit-based applications.
However, these works face key limitations:
  1. 1.
    Lack of High-Quality Data: Quantum programming lacks scalable, curated training datasets.
  2. 2.
    Code Complexity: Quantum circuits require LLMs to reason over non-trivial entanglement and measurement logic.
  3. 3.
    Hybrid Workflows: PennyLane introduces mixed classical-quantum code patterns that require deeper contextual understanding.
To bridge this gap, we propose the first PennyLane-centric dataset—PennyLang—paired with an advanced evaluation pipeline that incorporates RAG-based retrieval.
II-C Retrieval-Augmented Generation and GraphRAG
Retrieval-Augmented Generation (RAG) [47] has emerged as a popular method to ground LLM outputs using external context. In quantum programming, RAG helps models compensate for limited training data and outdated knowledge by retrieving relevant examples at inference time.
However, baseline RAG systems suffer from semantic drift, irrelevant context injection, and weak code relevance due to shallow vector-based similarity matching. To address this, we employ GraphRAG[32], a graph-enhanced retrieval framework designed to capture structural and functional relationships between quantum code samples. GraphRAG models PennyLang as a graph of interconnected examples, allowing retrievals that are topologically and semantically aligned with the user query.
Compared to traditional RAG, GraphRAG significantly improves context precision and reduces hallucinations in generated code. This leads to large gains in generation accuracy on the PennyLang test dataset.
Figure 2: Methodology for PennyLang quantum code generation and evaluation. Data is collected from GitHub, quantum books, and PennyLane docs, then refined, tokenized, and formatted for LLMs. The curated dataset is processed by GPT-4o for input-output conversion. User queries are handled via a retrieval-augmented generation (RAG) framework using LangChain and Chroma to fetch relevant code. Responses from GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B-Instruct-Turbo are compared against non-RAG outputs and evaluated on functionality, syntax, and modularity to assess RAG’s impact.
Our contributions advance AI-assisted quantum programming by enabling LLMs to provide high-quality, domain-specific assistance in PennyLane.
TABLE II: Comparison of Related Work and Our Contributions
Research Area Existing Works Our Contributions.
LLMs for Code Generation Codex [26], CodeBERT [43] PennyLane-specific dataset
LLMs for Quantum Programming Qiskit Assistant [22], Qiuchi et al. [46] Hybrid quantum-classical dataset
Quantum Code Datasets QCircuitNet [38], QDataset  [38] First PennyLane dataset
III PennyLang Framework
III-A Methodology for Data Collection and Refinement
Fig. 2 illustrates the methodology employed to construct a high-quality dataset of PennyLane code samples, ensuring that each snippet is genuinely relevant, well-structured, and accurately annotated. Our data collection strategy focused on extracting, filtering, and refining code snippets from multiple sources, including GitHub repositories, quantum computing books, and official PennyLane documentation.
III-A1 Data Collection from GitHub Repositories
To systematically gather PennyLane-specific code, we utilized the GitHub API to search for publicly accessible repositories with permissive open-source licenses (e.g., MIT, Apache 2.0, BSD). The selection criteria required that the term "PennyLane" be present in either the repository name, description, or README file.
To maintain dataset quality and avoid redundancy, we followed these constraints:
  • Only repositories from the main branch at the latest commit were considered.
  • Forked repositories were systematically excluded.
  • Only .py files (Python scripts) were extracted for further analysis.
  • Non-PennyLane files were filtered by checking for explicit import pennylane as qml statements.
  • Large, autogenerated files (e.g., test logs, package files) were excluded.
Since PennyLane maintains backward compatibility and rarely deprecates features [19], we did not impose a time constraint on data collection, ensuring a broad dataset scope. Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code.
This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [40].
III-A2 Extracting Quantum Code from Books
In addition to GitHub, we extracted PennyLane-related code snippets from two quantum computing books:
  • “Quantum Machine Learning: An Applied Approach” [48]
  • “A Practical Guide to Quantum Machine Learning and Quantum Optimization” [49]
These books provide theoretical foundations and applied examples of PennyLane-based quantum machine learning (QML) and variational quantum algorithms (VQAs). The extraction process involved:
  1. 1.
    Identifying all PennyLane-related code snippets.
  2. 2.
    Diversity Assurance: The dataset includes a broad spectrum of PennyLane use cases, from basic gate operations to advanced quantum optimization problems.
  3. 3.
    Extracting only the relevant code while discarding unrelated sections.
  4. 4.
    Manually verification of each snippet with explanations from the books, converting descriptive text into structured comments.
Unlike structured repositories, book content lacked a consistent format for where explanations were placed (before or after code). Thus, manual review was necessary to ensure that annotations correctly corresponded to each snippet.
III-A3 Scraping Code from PennyLane Documentation
The official PennyLane documentation [39] serves as a comprehensive reference for quantum programming, covering:
  • Quantum circuits and variational algorithms.
  • Quantum operators and templates.
  • Hybrid quantum-classical optimization.
  • Quantum chemistry applications.
TABLE III: Tokenized Input and Attention Masking.
Tokenized Input Attention Mask
"<pad>" 0
"<pad>" 0
"def" 1
":" 1
"circuit" 1
Tokenized Input Attention Mask
"(" 1
")" 1
"return" 1
"qml.expval" 1
To further enhance our dataset, we systematically scraped all code snippets from the official PennyLane website. This included examples embedded in tutorials and API references.
As with book sources, documentation explanations were manually converted into comments, ensuring that each code sample contained self-contained, contextualized information. This practice enhances LLM interpretability, as models trained on such data can generate both code and explanations.
III-A4 Dataset Summary and Statistics
The final dataset contains a total of 3,347 PennyLane code samples, as shown in Table IV.
TABLE IV: Dataset Composition.
Source Samples Collected
Unofficial GitHub repositories 1,952
GitHub (PennyLaneAI repository + others) 1,321
Official PennyLane Documentation 53
Quantum Computing Books 21
Total Dataset Size 3,347
Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for fine-tuning large language models (LLMs) for PennyLane-based quantum programming.
III-A5 Ensuring Data Integrity and Diversity
To maintain dataset integrity, we applied multiple quality control measures:
  1. 1.
    Duplicate Removal: Identical code samples from different sources were eliminated.
  2. 2.
    Code Formatting: Samples were formatted consistently to follow PEP 8 style guidelines [50].
  3. 3.
    Manual Review: A human-in-the-loop verification was conducted to assess annotation accuracy.
These efforts ensure that the dataset is rich, well-structured, and optimized for AI model training.
III-A6 Implications for AI-Driven PennyLane Code Generation
The collected dataset significantly enhances AI-driven PennyLane code assistance in multiple ways:
  • Improving Model Performance: High-quality annotations ensure that models understand both syntax and context.
  • Facilitating Explainable AI: Models trained on our dataset can generate explanatory comments alongside code.
This dataset serves as a foundation for future research in AI-assisted quantum programming for PennyLane.
TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework.
Component Function Description
DirectoryLoader Document Loader Loads PennyLane dataset samples as text files from a directory.
Chroma Vector Database Stores embeddings for retrieval using similarity search.
OpenAIEmbeddings Embedding Model Converts text into vector embeddings for indexing and retrieval.
ChatPromptTemplate Prompt Formatting Defines structured prompts for AI models to ensure consistency.
create_retrieval_chain Retrieval Pipeline Creates a pipeline that integrates stored embeddings for improved responses.
ChatOpenAI OpenAI LLM Interface Connects to OpenAI models (GPT-4o-mini) for response generation.
ChatAnthropic Anthropic LLM Interface Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses.
TABLE VI: LLMs Used in the RAG-Based Evaluation Framework.
Model Parameters Architecture Provider Usage in Pipeline
Qwen2.5-7B-Instruct-Turbo 7B Causal Decoder-Only Transformer Alibaba Cloud Generates PennyLane code
GPT-4o Mini Not disclosed Causal Decoder-Only Transformer OpenAI Evaluates and scores generated responses
Claude 3.5 Sonnet Not disclosed Causal Decoder-Only Transformer Anthropic Alternative model for response generation
III-B Tokenization and Formatting for LLM Training
This section details the preprocessing pipeline for preparing our dataset for (LLM) training/fine-tuning, focusing on tokenization, padding strategies, and attention masking to optimize input handling within the transformer architecture.
III-B1 Handling Variable-Length Code Samples
The collected PennyLane code samples exhibit varying lengths after tokenization. Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$. Following standard practice, left padding was applied to align shorter samples to $N$ tokens, ensuring batch processing efficiency and minimizing computational overhead.
III-B2 Padding Strategies and Tokenizer Variations
Padding strategies differ across tokenizers:
  • Token-Specific Padding: Some tokenizers have a dedicated padding token, automatically appended to shorter sequences. For example, Hugging Face’s Tokenizer uses "<pad>" by default [51].
  • Manually Defined Padding: If a tokenizer lacks an explicit padding token, a placeholder token (e.g., "0" or "<mask>") must be manually assigned.
  • Left vs. Right Padding: Left padding is commonly used in autoregressive models [52], ensuring that padding tokens do not interfere with causal attention mechanisms.
In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, and Claude 3.5 Sonnet. This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining autoregressive consistency.
III-B3 Applying Attention Masking
To improve transformer efficiency, an attention mask [53] was applied post-padding. This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data. Table III represents the attention mask matrix applied to a batch of tokenized samples. Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
  • Attention Mask Structure: A binary mask is assigned to each token:
    • 1 → Meaningful tokens to be processed by the model.
    • 0 → Padding tokens ignored by the transformer.
  • Impact on Training: The self-attention mechanism [52] ensures that padding tokens do not contribute to gradients, thereby reducing unnecessary computations and preventing misleading dependencies.
Figure 3: Distribution of instruction and response lengths with their correlation. The left plot shows the normal distribution of instruction lengths, the middle plot highlights the right-skewed distribution of response lengths, and the right scatter plot illustrates the relationship between instruction and response length.
III-C Methodology for Dataset Evaluation
III-C1 Retrieval-Augmented Generation (RAG) Framework
To assess the impact of context-aware prompting on LLM-based quantum code generation, we adopt a Retrieval-Augmented Generation (RAG) framework. This method enhances generation quality by incorporating relevant external context retrieved from our curated PennyLang dataset. RAG helps address common shortcomings in LLMs, such as hallucinations and outdated or incorrect syntax, by anchoring their output in concrete examples from real-world PennyLane programs. This results in improved syntactic validity, functional correctness, and alignment with PennyLane best practices.
III-C2 RAG Implementation using LangChain
Our RAG pipeline is implemented using the LangChain framework, which provides a modular infrastructure for document loading, vector embedding, retrieval, and prompt formatting. The dataset—formatted in an instruction-response structure—is first embedded using the OpenAI Embeddings module. The resulting vectors are stored in a Chroma database for efficient retrieval.
When a user query is submitted (e.g., a quantum programming task or challenge description), we apply Maximum Marginal Relevance (MMR) search to extract the top relevant examples from the dataset. These examples are concatenated into a structured prompt and fed into the LLM for code generation. To ensure fairness and evaluate RAG’s effectiveness, we generate responses both with and without retrieved context using the same query.
III-C3 RAG Evaluation
To benchmark the performance of RAG-augmented LLMs, we evaluate three state-of-the-art models: GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5-7B-Instruct-Turbo. Each model is tested on 29 quantum code generation tasks sampled from the QHack challenge dataset. For each query, both a baseline (non-RAG) and a RAG-enhanced response is generated. We then use GPT-4o to evaluate the outputs according to a rubric comprising three criteria:
  • Functionality: Does the code produce the correct quantum behavior?
  • Syntax: Is the code valid and executable in PennyLane?
  • Modularity: Is the solution well-organized and reusable?
Figure 4: Performance of GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5 with and without RAG. Mean scores (±1 SD) over 29 trials are shown for each model. Blue: RAG; Orange: Non-RAG.
Figure 4 summarizes the outcome of these evaluations. RAG-enhanced prompting consistently improves model performance across all criteria. On average, GPT-4o Mini improves by 11.67%, Claude 3.5 by 7.69%, and Qwen 2.5 by 14.38%. These gains highlight how retrieval provides critical context, especially for complex or multi-step quantum circuits.
III-C4 GraphRAG for Advanced Retrieval
To further improve retrieval precision beyond vector-based similarity, we extend our RAG framework using GraphRAG. While the baseline RAG implementation relies solely on Maximum Marginal Relevance (MMR) vector retrieval, GraphRAG incorporates structural relationships between code samples to guide context selection. In this method, the PennyLang dataset is organized into a graph where nodes represent code samples and edges represent functional or structural similarity—such as shared quantum gates, observables, and template usage patterns.
When a query is received, GraphRAG identifies a subgraph around the most semantically similar samples and performs hybrid retrieval by combining graph neighborhood traversal with traditional vector similarity. This enables the selection of more functionally aligned and topologically relevant code snippets, resulting in higher quality context for LLMs.
We evaluate GraphRAG using the same task set as our baseline experiments but restrict the evaluation to the GPT-4o Mini model due to its superior baseline performance. The model was tested on 146 PennyLang test tasks in both no-RAG and GraphRAG settings. The correctness of the generated outputs was determined by whether the solution passed the associated unit tests.
TABLE VII: Performance of GPT-4o Mini: No-RAG vs. GraphRAG (146 tasks).
Metric No-RAG GraphRAG
Correct Solutions 30 85
Incorrect Solutions 116 61
Total Samples 146 146
Accuracy (%) 20.55% 58.22%
As shown in Table VII, the GraphRAG approach yields a dramatic improvement in accuracy—from 20.55% without retrieval to 58.22% with graph-enhanced retrieval. This threefold increase highlights the value of graph-based structural reasoning in quantum programming tasks, particularly for problems involving complex circuit construction or uncommon PennyLane APIs.
Together, these results establish GraphRAG as a significant enhancement to conventional RAG pipelines and demonstrate its practical utility in improving the performance of LLMs on challenging real-world quantum programming benchmarks.
IV Statistical Evaluation of the Dataset
A comprehensive analysis of the dataset was conducted to assess the distribution of instruction lengths, response lengths, response variations across feature groups, and their correlation with quantum functionalities. This section provides insights into the dataset’s structure, response variability, and its implications for training LLMs.
IV-A Response Lengths Across Feature Groups
Figure 5 provides a box plot representation of response lengths across different feature groups.
  • Response lengths remain relatively stable across feature groups, suggesting uniformity in generated outputs.
  • Outliers appear in every feature category, indicating occasional long responses, particularly in core operations, measurement, and mathematical functions.
  • The “Other” category has the shortest responses, likely due to simpler queries requiring minimal output.
Figure 5: Box plot of response lengths across different feature groups. Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions.
IV-B Instruction and Response Length Distributions
Figure 3 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset. Instruction lengths, as shown in the left plot, follow a normal distribution with a peak around 35 words. The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare. In contrast, response lengths exhibit a right-skewed distribution, as shown in the middle plot. Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs. The instruction-response correlation, illustrated in the rightmost scatter plot, reveals a positive correlation between instruction and response length. While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
IV-C Feature Group Usage Across Quantum Domains
The relationship between feature groups and quantum computing domains, such as machine learning, hardware, chemistry, and optimization, is illustrated in Figure 6. Quantum Machine Learning relies predominantly on quantum gates and operators, alongside error handling techniques, reflecting its dependence on structured quantum transformations and noise mitigation. In contrast, Quantum Chemistry and Optimization exhibit relatively lower usage of mathematical functions, suggesting that these domains rely on alternative computational approaches rather than extensive arithmetic operations. Quantum Hardware applications, however, demonstrate a balanced distribution across various feature groups, indicating a broader, more integrated usage of quantum functionalities across different operational layers.
Figure 6: Feature Group Usage Across Domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories, with color intensity reflecting count.
IV-D Feature Category Association with Quantum Functions
To analyze how different quantum functionalities relate to feature categories, we examine the confusion matrix in Figure 7.
Figure 7: Confusion matrix showing the relationship between newly defined semantic categories and various PennyLane quantum features. The color intensity represents the frequency of occurrences, where darker shades indicate higher associations between a given category (rows) and a specific quantum feature (columns). This visualization provides insights into how different quantum functionalities are distributed across conceptual categories, aiding dataset structuring and LLM-based learning tasks.
Quantum gates and operators exhibit the strongest associations with multiple features, particularly with commonly used functions such as qml.PauliZ, qml.RX, and qml.RY, which are integral to quantum circuit construction. Core operations and measurement functions also display high correlations, reflecting their frequent application in quantum circuit execution and result interpretation. Meanwhile, error handling and mathematical functions show relatively lower but still significant interactions with select quantum features, indicating their importance in specific computational scenarios rather than across all use cases.
IV-E Insights and Implications
The findings from Figures 3, 5, 7, and 6 offer several key insights:
  • Balanced Instruction Lengths: The normal distribution of instruction lengths ensures the dataset is well-structured for model training.
  • Response Length Variability: Most responses are concise, but outliers indicate that certain quantum features demand more detailed explanations.
  • Feature-Specific Trends: Quantum gates and measurement functions dominate dataset interactions, aligning with core PennyLane applications.
  • Domain-Specific Usage: Different quantum fields rely on distinct feature sets, highlighting potential customization opportunities for specialized LLM training.
V Case Studies
The case studies compare base LLM-generated PennyLane code with RAG-enhanced implementations, showcasing notable improvements in functionality, syntax, and modularity across different quantum programming tasks.
In the first task, implementing a 1-qubit parameterized RX gate, the base model introduced redundant imports, suboptimal measurement choices, and missing hardware-specific optimizations. The RAG-optimized implementation addressed these issues by incorporating hardware-aware device initialization, optimized measurement practices, and type annotations, achieving a perfect 10/10 functionality score.
The second task, developing an RX gate circuit with optional parameter binding, revealed static parameter handling and the absence of deferred execution in the base LLM output. The RAG-enhanced version introduced flexible parameter binding, probabilistic measurement strategies, and deferred execution patterns, improving adaptability and modularity.
The third case study, constructing a variational quantum classifier using StronglyEntanglingLayers, further demonstrated the superiority of RAG-based generation. The base LLM solution lacked adaptive training integration, efficient optimization, and TensorFlow-Keras support, requiring manual gradient updates. In contrast, the RAG-optimized model seamlessly integrated TensorFlow, Keras, and the Adam optimizer, significantly enhancing training efficiency and parameter optimization.
Overall, across all tasks, RAG-assisted code generation consistently outperformed the base LLM, producing more structured, modular, and functionally superior PennyLane code. This highlights the effectiveness of retrieval-augmented approaches in enhancing quantum programming workflows.
Task: Implement 1-qubit parameterized RX gate using PennyLane
Base LLM Implementation (Without RAG) Limitations Identified: Unnecessary numpy import Suboptimal measurement choice (expectation value vs state) Missing hardware-specific optimizations 1import pennylane as qml 2from pennylane import numpy as np # Redundant import 3 4dev = qml.device(’default.qubit’, wires=1) 5 6@qml.qnode(dev) 7def circuit(theta): 8 qml.RX(theta, wires=0) 9 return qml.expval(qml.PauliZ(0)) # Non-optimal measurement Listing 1: Non-RAG Implementation Evaluation Results: Functionality: 9/10 (Valid but limited measurement) Syntax: 9/10 (Redundant imports) Modularity: 8/10 (Missing type hints)
RAG-Enhanced Implementation Retrieved Context: Hardware-aware device initialization (sample_1137.txt) State measurement best practices (sample_214.txt) Testing methodologies (sample_23.txt) 1import pennylane as qml 2 3dev = qml.device("lightning.qubit", wires=1) # Hardware-optimized 4 5@qml.qnode(dev) 6def circuit(theta: float): # Type hinting 7 qml.RX(theta, wires=0) 8 return qml.state() # Full state measurement Listing 2: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Complete state analysis) Syntax: 10/10 (Clean imports) Modularity: 9/10 (Type annotations)
Task: Create 1-qubit RX gate circuit with optional parameter binding
Base LLM Implementation (Without RAG) Limitations Identified: Static parameter handling Missing deferred parameterization Limited measurement flexibility 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 if theta: # Binary parameter check 6 qml.RX(theta, wires=0) 7 return qml.expval(qml.PauliZ(0)) # Fixed measurement Listing 3: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (No variable binding) Syntax: 9/10 (Valid but rigid) Modularity: 8/10 (No late binding support)
RAG-Enhanced Implementation Retrieved Context: Variable parameter techniques (sample_1137.txt) Probabilistic measurement strategies (sample_214.txt) Deferred execution patterns (sample_23.txt) 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 param = theta if theta else qml.variable(’theta’) 6 qml.RX(param, wires=0) 7 return qml.probs(wires=0) # Full probability measurement Listing 4: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Flexible parameter binding) Syntax: 9/10 (Clear variable handling) Modularity: 9/10 (Supports late binding)
Task: Implement a Variational Quantum Classifier with Strongly Entangling Layers
Base LLM Implementation (Without RAG) Limitations Identified: Uses a static quantum circuit with fixed parameters, lacking adaptive training integration. Lacks integration with TensorFlow/Keras, requiring a manual optimization loop. Uses Gradient Descent Optimizer instead of the Adam optimizer, which may slow convergence. 1import pennylane as qml 2import numpy as np 3from sklearn.datasets import make_classification 4from sklearn.model_selection import train_test_split 5from sklearn.preprocessing import StandardScaler 6from sklearn.metrics import accuracy_score 7 8# Generate a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Scale the features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define the quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Define the quantum circuit 21@qml.qnode(dev) 22def circuit(params, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.StronglyEntanglingLayers(params, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 5: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (Works but lacks adaptive training integration) Syntax: 9/10 (Code is valid but manual optimization is not ideal) Modularity: 8/10 (Training and inference should be better structured)
RAG-Enhanced Implementation Retrieved Context: TensorFlow-Keras integration for efficient optimization. Adam optimizer instead of manual updates for improved convergence. Improved quantum circuit design based on best practices. 1import pennylane as qml 2import numpy as np 3import tensorflow as tf 4from sklearn.datasets import make_classification 5from sklearn.model_selection import train_test_split 6from sklearn.preprocessing import StandardScaler 7 8# Create a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Standardize features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Quantum circuit with Strongly Entangling Layers 21@qml.qnode(dev) 22def quantum_circuit(weights, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.templates.StronglyEntanglingLayers(weights, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 6: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Seamless Keras integration for training) Syntax: 10/10 (Concise and efficient TensorFlow pipeline) Modularity: 9/10 (Well-structured training process)
VI Conclusion
In this paper, we introduce PennyLang, a high-quality, instruction-response dataset tailored for PennyLane-based quantum code generation. We propose a robust data collection and annotation methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples. To assess the practical utility of the dataset, we develop a Retrieval-Augmented Generation (RAG) evaluation framework and benchmark three state-of-the-art LLMs—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B—under both RAG and non-RAG conditions. Our findings demonstrate consistent performance gains with RAG, especially in functionality and syntax correctness. Building on this, we introduce a GraphRAG extension that models structural relationships across code snippets using graph-based retrieval. When applied to GPT-4o Mini, GraphRAG yields a 2.8$\times$ improvement in accuracy compared to the non-RAG baseline, highlighting its potential for deeper semantic retrieval in quantum programming. We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development. Future directions include training lightweight instruction-tuned models on this dataset, scaling GraphRAG across different models and quantum frameworks, and expanding toward cross-framework generalization between PennyLane, Qiskit, and Cirq.
Acknowledgment
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008, and Center for CyberSecurity (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104. This research was carried out on the High Performance Computing resources at New York University Abu Dhabi.