### v3

# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

Abstract
Large Language Models (LLMs) offer powerful capabilities in code generation, natural language understanding, and domain-specific reasoning. Their application to quantum software development remains limited, in part because of the lack of high-quality datasets both for LLM training and as dependable knowledge sources. To bridge this gap, we introduce PennyLang, an off-the-shelf, high-quality dataset of 3,347 PennyLane-specific quantum code samples with contextual descriptions, curated from textbooks, official documentation, and open-source repositories. Our contributions are threefold: (1) the creation and open-source release of PennyLang, a purpose-built dataset for quantum programming with PennyLane; (2) a framework for automated quantum code dataset construction that systematizes curation, annotation, and formatting to maximize downstream LLM usability; and (3) a baseline evaluation of the dataset across multiple open-source models, including ablation studies, all conducted within a retrieval-augmented generation (RAG) pipeline. Using PennyLang with RAG substantially improves performance: for example, Qwen 7B’s success rate rises from 8.7% without retrieval to 41.7% with full-context augmentation, and LLaMa 4 improves from 78.8% to 84.8%, while also reducing hallucinations and enhancing quantum code correctness. Moving beyond Qiskit-focused studies, we bring LLM-based tools and reproducible methods to PennyLane for advancing AI-assisted quantum development.
Index Terms:
Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), GraphRAG, Code Completion and Assistance, Dataset Curation, Evaluation Framework
I Introduction
Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [1, 2, 3]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [4], where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer, highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [17, 18]. PennyLane is a leading open-source Python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [19]. However, unlike the Qiskit framework [20, 21], which benefits from dedicated AI-driven code assistants [22], PennyLane lacks equivalent tools to assist developers in writing optimized quantum code.
In parallel, the field of Natural Language Processing (NLP) has seen rapid advancements in Large Language Models (LLMs), which now excel at reasoning, contextual understanding, and automated code generation [23, 24, 25]. These models have shown success in classical programming [26, 27], improving developer productivity and reducing syntax and logic errors. However, their application in quantum programming, especially within the PennyLane ecosystem, remains underexplored. This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for refinement, annotation, and evaluation, ultimately enabling LLM-based quantum code generation and debugging.
State-of-the-Art and Limitations: LLMs have been extensively benchmarked for classical programming [28, 29], but their performance in quantum domains remains less explored. Early efforts such as Qiskit HumanEval and Qiskit Code Assistant [22] adapt LLMs to quantum programming but focus on Qiskit’s hardware-centric paradigm. PennyLane, by contrast, targets variational quantum algorithms and quantum machine learning [30], requiring dedicated LLM support and a specialized training corpus. A major barrier is the scarcity of suitable data: quantum code is scattered across papers, forums, repositories, and documentation [31], and often lacks the contextual annotations necessary for LLMs to link quantum operations with their intended use.
Figure 1: Overview of novel contributions in this work.
Our Approach and Contributions: In this work, we introduce PennyLang, a large-scale, high-quality dataset specifically designed for PennyLane code generation, accompanied by a rigorous evaluation pipeline. We implement a baseline Retrieval-Augmented Generation (RAG) framework to assess the dataset’s effectiveness across multiple models. Our key contributions, shown in Fig. 1, consist of:
  1. 1.
    Comprehensive PennyLane Dataset: Curated 3,347 PennyLane-specific code samples from GitHub repositories, quantum computing textbooks, and official PennyLane documentation, each enriched with contextual descriptions and annotations to support instruction tuning.
  2. 2.
    Systematic Data Refinement and Annotation: Developed a robust pipeline for dataset cleaning, metadata enrichment, and conversion of code snippets into instruction–query pairs using the GPT-4o API, ensuring compatibility with LLM training workflows.
  3. 3.
    RAG-Based Evaluation Framework: Designed a RAG evaluation framework to assess the dataset’s impact on model performance. Evaluated GPT-4o mini, Claude 3.5, Qwen 7B, and LLaMa 4 Maverick 17B under both retrieval-augmented and non-augmented settings, scoring outputs on functional correctness.
Summary of Key Results: Retrieval augmentation with PennyLang yields substantial relative gains for smaller open-source models: Qwen 7B’s success rate rises from 8.71% to 41.66% ($\approx$32.95% relative increase), and LLaMa 4 improves from 78.78% to 84.84% ($\approx$7.7% relative increase). In contrast, stronger baselines (GPT-4o-mini and Claude 3.5 Sonnet) do not benefit from naive full-context retrieval, their best performance is achieved without retrieval (85.60% and 95.07%, respectively), with full retrieval resulting in slight degradation. Detailed analysis is presented in Section 5.
II Background and Related Work
TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training
Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Framework Datasets Qiskit Textbook Dataset [32] Qiskit Qiskit.org Textbook Multiple Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [33] Qiskit IBM Quantum Experience Derived Instruction-Response Format JSON/Notebook Qiskit Code Repository [34] Qiskit GitHub (Qiskit AI repos) Thousands Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [35] Qiskit Kaggle (Possible) 309 Pretrained Benchmark Data CSV/JSON QDataSet [36] Qiskit GitHub 52 Quantum ML Data Various QCircuitNet [37] Qiskit Research Paper 4,800 Quantum Algorithm Design Benchmark Python/QASM PennyLane Framework Datasets PennyLane Documentation [38] PennyLane PennyLane.ai Docs Various Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [39] PennyLane GitHub Open-Source Scattered Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Curated and Structured Dataset JSON file
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [40] and automated code generation [41], their application in quantum software development remains underexplored. Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [22], with little attention to the PennyLane framework [19], despite its popularity in hybrid quantum-classical workflows. This section presents an in-depth review of related work in three key areas:
  1. 1.
    LLMs for Code Generation in Classical Programming
  2. 2.
    LLMs for Quantum Programming
  3. 3.
    Datasets for Training LLMs in Quantum Code Generation (summarized in Table I)
We then summarize existing contributions and highlight our novel approach in developing a PennyLane-centric dataset to advance AI-driven quantum programming, as shown in Table II.
II-A LLMs for Code Generation in Classical Programming
Transformer-based language models have revolutionized automated code generation, significantly advancing classical software development. OpenAI’s Codex [26], trained on large-scale public code repositories, enabled natural language-to-code translation, setting a new standard for programming assistants. CodeBERT [42] further improved source code understanding through bidirectional transformers, while InCoder [29] introduced multi-turn interactive capabilities for code infilling and synthesis. AlphaCode [25] demonstrated that LLMs could generate solutions for competitive programming problems, reaching near-human performance. To rigorously evaluate these models, benchmarks such as HumanEval [24], MBPP [43], and CodeXGLUE [44] have been proposed, each offering complementary perspectives on correctness, diversity, and execution fidelity. These benchmarks collectively emphasize the importance of well-curated, structured datasets in improving LLM-driven code generation.
TABLE II: Comparison: Related Work & Our Contributions
Research Area Existing Works Our Contributions LLMs for Code Generation Codex [26] CodeBERT [42] PennyLane-specific dataset LLMs for Quantum Programming Qiskit Assistant [22] [45] Hybrid quantum- classical dataset Quantum Code Datasets QCircuitNet [37] QDataset [37] First PennyLane dataset
II-B LLMs for Quantum Programming
In comparison, quantum programming has seen limited adoption of LLMs, with most efforts focused on the Qiskit framework. IBM’s Qiskit Code Assistant [22] supports Qiskit programs’ debugging and optimizations, while [45] adapted BERT-based models for enhanced quantum code comprehension. CodeLlama-Quantum [27] extended Meta’s CodeLlama to address Qiskit-specific use cases. Despite these advances, key limitations persist. The field lacks large, high-quality datasets for training and evaluation, and quantum circuits demand deeper reasoning due to entanglement, measurement, and complex control logic. Furthermore, hybrid programming paradigms like PennyLane interleave classical and quantum instructions, requiring models to understand multi-modal code patterns and contextual dependencies. To address these challenges, we propose PennyLang, the first PennyLane-centric dataset, supported by an evaluation pipeline that integrates RAG to improve code relevance and contextual grounding. Our contributions advance AI-assisted quantum programming by enabling LLMs to provide high-quality, domain-specific assistance in PennyLane.
Figure 2: Methodology for PennyLang Quantum Code Generation and Evaluation. We collect and refine quantum code from GitHub, textbooks, and PennyLane documentation, then use GPT-4o to convert cleaned snippets into instruction–query pairs (with code and tests). These examples are embedded in Chroma, and LangChain performs MMR-based retrieval to produce RAG-augmented versus vanilla prompts. GPT-4o-mini, Claude 3.5 Sonnet, LLaMa 4-17B Maverick, and Qwen 2.5-7B-Instruct-Turbo are evaluated on functional correctness to quantify the benefit of PennyLang dataset.
III PennyLang Framework
III-A Methodology for Data Collection and Refinement
Fig. 2 illustrates the methodology employed to construct a high-quality dataset of PennyLane code samples, ensuring that each snippet is genuinely relevant, well-structured, and accurately annotated. Our data collection strategy focused on extracting, filtering, and refining code snippets from multiple sources, including GitHub repositories, quantum computing books, and official PennyLane documentation.
III-A1 Data Collection from GitHub Repositories
To systematically gather PennyLane-specific code, we utilized the GitHub API to search for publicly accessible repositories with permissive open-source licenses (e.g., MIT, Apache 2.0, BSD). The selection criteria required that the term "PennyLane" be present in either the repository name, description, or README file. To maintain dataset quality and avoid redundancy, we followed these constraints:
  • Only repositories from the main branch at the latest commit were considered.
  • Forked repositories were systematically excluded.
  • Only Python (.py) files were extracted for further analysis.
  • Non-PennyLane files were filtered by checking for explicit import pennylane as qml statements.
  • Large, autogenerated files (e.g., test logs, package files) were excluded.
Since PennyLane maintains backward compatibility and rarely deprecates features [19], we did not impose a time constraint on data collection, ensuring a broad dataset scope. Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code. This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [39].
III-A2 Extracting Quantum Code from Books
In addition to GitHub, we extracted PennyLane-related code snippets from two quantum computing books, Quantum Machine Learning: An Applied Approach  [46] and A Practical Guide to Quantum Machine Learning and Quantum Optimization [47]. These books provide theoretical foundations and applied examples of PennyLane-based quantum machine learning (QML) and variational quantum algorithms (VQAs). The extraction process involved:
  1. 1.
    Identifying all PennyLane-related code snippets.
  2. 2.
    Diversity Assurance: The dataset includes a broad spectrum of PennyLane use cases, from basic gate operations to advanced quantum optimization problems.
  3. 3.
    Extracting only the relevant code while discarding unrelated sections.
  4. 4.
    Manually verification with explanations from the books, converting descriptive text into structured comments.
Unlike structured repositories, book content lacked a consistent format for where explanations were placed (before or after code). Thus, manual review was necessary to ensure that annotations correctly corresponded to each snippet.
III-A3 Scraping Code from PennyLane Documentation
The official PennyLane documentation [38] is a comprehensive resource covering topics such as quantum circuits, variational algorithms, operators and templates, hybrid quantum–classical optimization, and quantum chemistry applications. To enrich our dataset, we systematically extracted all code snippets from the official website, including those embedded in tutorials and API references. Documentation text accompanying each example was manually converted into inline comments, ensuring every sample is self-contained and context-rich. This approach improves LLM interpretability, enabling models trained on the dataset to produce both functional code and explanatory text.
III-A4 Dataset Summary and Statistics
The final dataset containing 3,347 PennyLane code samples is shown in Table III.
TABLE III: Dataset Composition
Source Samples Collected Unofficial GitHub repositories 1,952 GitHub (PennyLaneAI repository + others) 1,321 Official PennyLane Documentation 53 Quantum Computing Books 21 Total Dataset Size 3,347
Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for enhancing LLMs for PennyLane-based quantum programming.
III-A5 Ensuring Data Integrity and Diversity
To keep dataset integrity, we applied multiple quality control measures:
  1. 1.
    Duplicate Removal: Identical code samples from different sources were eliminated.
  2. 2.
    Code Formatting: Samples were formatted consistently to follow PEP 8 style guidelines [48].
  3. 3.
    Manual Review: A human-in-the-loop verification was conducted to assess annotation accuracy.
These efforts ensure that the dataset is rich, well-structured, and optimized for AI model training.
III-A6 Implications for AI-Driven PennyLane Code Generation
The collected dataset significantly enhances AI-driven PennyLane code assistance in multiple ways:
  • Improving Model Performance: High-quality annotations ensure models understand both syntax and context.
  • Facilitating Explainable AI: Models trained on PennyLang can generate explanatory comments alongside code.
This dataset serves as a foundation for future research in AI-assisted quantum programming for PennyLane.
III-B Tokenization and Formatting for LLM Training
This section details the preprocessing pipeline for preparing our dataset for (LLM) training/fine-tuning, focusing on tokenization, padding strategies, and attention masking to optimize input handling within the transformer architecture.
III-B1 Handling Variable-Length Code Samples
The collected PennyLane code exhibit varying lengths after tokenization. Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$. Following standard practice, left padding was applied to align shorter samples to $N$ tokens, ensuring batch processing efficiency and minimizing computational overhead.
III-B2 Padding Strategies and Tokenizer Variations
Padding strategies differ across tokenizers:
  • Token-Specific Padding: Some tokenizers have a dedicated padding token, automatically appended to shorter sequences. For example, Hugging Face’s Tokenizer uses "<pad>" by default [49].
  • Manually Defined Padding: If a tokenizer lacks an explicit padding token, a placeholder token (e.g., "0" or "<mask>") must be manually assigned.
  • Left vs. Right Padding: Left padding is commonly used in autoregressive models [50], ensuring that padding tokens do not interfere with causal attention.
In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, Claude 3.5 Sonnet, and LLaMa 4. This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining auto-regressive consistency.
III-B3 Applying Attention Masking
To improve transformer efficiency, an attention mask [51] was applied post-padding. This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data. Table IV shows the attention mask matrix applied to a batch of tokenized samples. Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
TABLE IV: Tokenized Input and Attention Masking.
Token Mask Token Mask Token Mask "<pad>" 0 "def" 1 "circuit" 1 "<pad>" 0 ":" 1 "(" 1 ")" 1 "return" 1 "qml.expval" 1
  • Attention Mask Structure: A binary mask is assigned to each token:
    • 1 → Meaningful tokens to be processed by the model.
    • 0 → Padding tokens ignored by the transformer.
  • Impact on Training: The self-attention mechanism [50] ensures that padding tokens do not contribute to gradients, thereby reducing unnecessary computations and preventing misleading dependencies.
III-C Methodology for Dataset Evaluation
III-C1 Retrieval-Augmented Generation (RAG) Framework
To improve PennyLane code generation by reducing hallucinations and outdated syntax—and to better align outputs with PennyLane best practices—we adopt a retrieval-augmented generation (RAG) [52] approach. We implement the pipeline in LangChain with modular components for document ingestion, vector embedding, retrieval, and prompt construction. The PennyLang dataset, formatted as instruction–response pairs, is embedded using OpenAIEmbeddings and stored in a Chroma vector database for efficient similarity search. At inference, user queries are embedded, relevant context is retrieved, and the combined prompt is sent to the target LLM. Table V summarizes the key PennyLang RAG framework components, including their role in dataset loading, embedding storage, prompt formatting, retrieval chaining, and interfacing with multiple LLM backends.
TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework
PennyLang RAG Framework Components DirectoryLoader (Document Loader)    $\triangleright$ Loads PennyLane dataset samples as text files from directory Chroma (Vector Database)    $\triangleright$ Stores embeddings for retrieval using similarity search OpenAIEmbeddings (Embedding Model)    $\triangleright$ Converts text into vector embeddings for indexing and retrieval ChatPromptTemplate (Prompt Formatting)    $\triangleright$ Defines structured prompts for AI models to ensure consistency create_retrieval_chain (Retrieval Pipeline)    $\triangleright$ Creates pipeline integrating stored embeddings for improved responses ChatOpenAI (OpenAI LLM Interface)    $\triangleright$ Connects to OpenAI models (GPT-4o-mini) for response generation ChatAnthropic (Anthropic LLM Interface)    $\triangleright$ Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses
When a user query is submitted (e.g., a quantum programming task or challenge description), we use Maximum Marginal Relevance (MMR) search to retrieve the most relevant examples from the dataset. The retrieved examples are concatenated into a structured prompt and passed to the LLM for code generation. For a fair assessment of RAG’s impact, we generate outputs for each query both with and without retrieved context under identical conditions.
IV Statistical Evaluation of the Dataset
A comprehensive analysis of the dataset was conducted to assess the distribution of instruction lengths, response lengths, response variations across feature groups, and their correlation with quantum functionalities. This section provides insights into the dataset’s structure, response variability, and its implications for training LLMs.
IV-A Response Lengths Across Feature Groups
Figure 3 provides a box plot representation of response lengths across different feature groups.
  • Response lengths remain relatively stable across feature groups, suggesting uniformity in generated outputs.
  • Outliers appear in every feature category, indicating occasional long responses, particularly in core operations, measurement, and mathematical functions.
  • The “Other” category has the shortest responses, likely due to simpler queries requiring minimal output.
Figure 3: Box plot of response lengths across different feature groups. Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions.
Figure 4: Distribution of instruction and response lengths with their correlation. The left plot shows the normal distribution of instruction lengths, the middle plot highlights the right-skewed distribution of response lengths, and the right scatter plot illustrates the relationship between instruction and response length.
IV-B Instruction and Response Length Distributions
Figure 4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset. Instruction lengths, as shown in left plot, follow a normal distribution with a peak around 35 words. The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare. In contrast, response lengths exhibit a right-skewed distribution, as shown in the middle plot. Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs. The instruction-response correlation, illustrated in the rightmost plot, reveals a positive correlation between instruction and response length. While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
IV-C Feature Group Usage Across Quantum Domains
The relationship between feature groups and quantum computing domains, such as machine learning, hardware, chemistry, and optimization, is illustrated in Figure 5. Quantum Machine Learning relies predominantly on quantum gates and operators, alongside error handling techniques, reflecting its dependence on structured quantum transformations and noise mitigation. In contrast, Quantum Chemistry and Optimization exhibit relatively lower usage of mathematical functions, suggesting that these domains rely on alternative computational approaches rather than extensive arithmetic operations. Quantum Hardware applications, however, demonstrate a balanced distribution across various feature groups, indicating a broader, more integrated usage of quantum functionalities across different operational layers.
Figure 5: Feature group usage across domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories (QML, Quantum Hardware, Quantum Chemistry and Optimization), with color intensity reflecting count.
TABLE VI: Evaluation of Qwen 2.5 7B, LLaMa 4 Maverick 17B, GPT-4o-mini, and Claude 3.5 Sonnet on PennyLane
Metric Qwen 7B LLaMa 4 GPT-4o-mini Claude 3.5 Sonnet Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Total Notebooks 1320 1320 1320 1320 Working Notebooks 467 495 480 95 901 875 895 799 712 780 755 625 899 911 856 1001 Success Rate (%) 41.66 45.07 43.18 8.71 84.84 84.84 82.17 78.78 80.30 84.84 81.81 85.60 89.01 92.44 87.50 95.07 Pass@1 87 100 88 19 179 171 179 150 144 158 149 134 187 182 170 194 Pass@2 99 109 103 21 198 202 199 178 171 189 175 175 210 218 203 230 Pass@3 104 113 109 21 212 214 209 191 190 207 195 199 223 232 214 238 Pass@4 109 117 110 22 216 217 212 202 202 219 209 215 230 242 230 246 Pass@5 110 119 114 23 224 224 217 208 212 224 216 226 235 244 231 251
IV-D Feature Category Association with Quantum Functions
To analyze how different quantum functionalities relate to feature categories, we examine the confusion matrix in Figure 6.
Figure 6: Confusion matrix showing the relationship between newly defined semantic categories and various PennyLane quantum features. The color intensity represents the frequency of occurrences, where darker shades indicate higher associations between a given category (rows) and a specific quantum feature (columns). This visualization provides insights into how different quantum functionalities are distributed across conceptual categories, aiding dataset structuring and LLM-based learning tasks.
Quantum gates and operators exhibit the strongest associations with multiple features, particularly with commonly used functions such as qml.PauliZ, qml.RX, and qml.RY, which are integral to quantum circuit construction. Core operations and measurement functions also display high correlations, reflecting their frequent application in quantum circuit execution and result interpretation. Meanwhile, error handling and mathematical functions show relatively lower but still significant interactions with select quantum features, indicating their importance in specific computational scenarios rather than across all use cases.
IV-E Insights and Implications
Findings from Figures 4, 3, 6, and 5 offer several insights:
  • Balanced Instruction Lengths: The normal distribution of instruction lengths ensures the dataset is well-structured for model training.
  • Response Length Variability: Most responses are concise, but outliers indicate that certain quantum features demand more detailed explanations.
  • Feature-Specific Trends: Quantum gates and measurement functions dominate dataset interactions, aligning with core PennyLane applications.
  • Domain-Specific Usage: Different quantum fields rely on distinct feature sets, highlighting potential customization opportunities for specialized LLM training.
V Evaluation
To assess the effectiveness of our proposed dataset for quantum code generation, we designed a benchmark suite comprising 264 test cases, each tailored to evaluate various PennyLane functionalities and quantum programming concepts. Each test was used as a prompt within an RAG pipeline and evaluated across both open-source and commercial LLMs. The evaluation utilized the Pass@5 metric, where each model generates five candidate solutions per test, and success rate (%) is computed as $\text{pass@5}/{264}\times 100$, where pass@5 indicates the number of test cases for which at least one correct solution was generated among the five attempts. In total, 1320 notebooks were generated and executed per model (264 tests × 5 completions), ensuring a rigorous, code-level assessment of correctness. Table VII lists the details for evaluated models.
TABLE VII: LLMs Used in the RAG-Based Evaluation Framework
Large Language Models in RAG Pipeline Qwen2.5-7B-Instruct-Turbo (Alibaba Cloud)    $\bullet$ Parameters: 7B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Generates PennyLane code LLaMa 4 Maverick (Meta)    $\bullet$ Parameters: 17B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Code generation (planned evaluation) GPT-4o Mini (OpenAI)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Evaluates and scores generated responses Claude 3.5 Sonnet (Anthropic)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Alternative model for response generation
Open-Source Models.
Given that the dataset was primarily developed to support open-source development, our core evaluation focused on Qwen 7B and LLaMa 4. These models showed clear improvements when used with our dataset. Notably, RAG with 75% context coverage often outperformed the full-context setting. This suggests that a moderate amount of relevant context may reduce noise and improve grounding, while excessive retrieval might introduce redundancy or irrelevant tokens.
Commercial Models.
For comparison, we also tested GPT-4o-mini and Claude 3.5. As expected, these commercial models demonstrated high performance even without RAG. However, the added context from our dataset provided limited benefit. This may be attributed to their broad pretraining, which likely includes similar material (e.g., PennyLane documentation or equivalent quantum code examples), reducing the marginal gain from external retrieval.
Discussion.
The evaluation reveals important interactions between dataset design, model capacity, and retrieval strategies. Open-source models such as Qwen 7B and LLaMa 4 showed notable improvements when augmented with PennyLang, confirming that domain-specific RAG can effectively compensate for limited pretraining in quantum programming. Interestingly, the 75% context setting consistently outperformed the full-context retrieval, indicating that focused and moderately sized context windows improve model grounding by avoiding redundancy and context saturation.
In contrast, commercial models like GPT-4o Mini and Claude 3.5 achieved high performance across all settings, with only marginal benefits from added retrieval. This behavior is likely due to their extensive pretraining on diverse codebases, which may already include material similar to what our dataset provides. Consequently, the impact of external retrieval is diminished and can even introduce noise if not precisely targeted. These findings emphasize the importance of retrieval quality over quantity in RAG-based code generation. The consistent advantage of partial over full context suggests that carefully curated, relevant examples are more beneficial than exhaustive inclusion.
Finally, by evaluating 1,320 generated notebooks per model, we ensure that our conclusions are grounded in executable and functionally correct outputs, a crucial criterion in the domain of quantum programming. These results underscore the utility of PennyLang in enhancing open-source LLM performance while providing insights into optimal RAG configurations for technical code generation tasks.
VI Conclusion
In this paper, we introduce PennyLang, a high-quality instruction–response dataset tailored for PennyLane-based quantum code generation. We present our data collection and prepossessing methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples. To evaluate its practical utility, we develop a RAG evaluation framework and benchmark four state-of-the-art LLMs (GPT-4o Mini, Claude 3.5 Sonnet, Qwen2.5-7B, and LLaMa 4-17B Maverick) under both RAG and non-RAG settings. Our findings show that PennyLang, when paired with RAG, substantially boosts performance for open-source models, while proprietary models exhibit minimal or negative gains with naive full-context retrieval, likely because their pre-training already covers overlapping PennyLane-related knowledge. This further highlights the importance of calibrated retrieval budgets, as moderate settings (e.g., 75%) often outperform full context by balancing signal and noise. We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development. Future directions include training lightweight instruction-tuned models on this dataset, optimizing retrieval strategies across different model types, and extending to cross-framework generalization between PennyLane, Qiskit, and Cirq.
Acknowledgment
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008, and Center for CyberSecurity (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104. This research was carried out on the High Performance Computing resources at New York University Abu Dhabi.